### v1

# RiTeK: A Dataset for Large Language Models Complex Reasoning over Textual Knowledge Graphs

Abstract
Answering complex real-world questions often requires accurate retrieval from textual knowledge graphs (TKGs). The scarcity of annotated data, along with intricate topological structures, makes this task particularly challenging. As the nature of relational path information could enhance the inference ability of Large Language Models (LLMs), efficiently retrieving more complex relational path information from TKGs presents another key challenge. To tackle these challenges, we first develop a Dataset for LLMs Complex Reasoning over Textual Knowledge Graphs (RiTeK) with a broad topological structure coverage. We synthesize realistic user queries that integrate diverse topological structures, relational information, and complex textual descriptions. We conduct rigorous expert evaluation to validate the quality of our synthesized queries. And then, we introduce an enhanced Monte Carlo Tree Search (MCTS) method, Relational MCTS, to automatically extract relational path information from textual graphs for specific queries. Our dataset mainly covers the medical domain as the relation types and entity are complex and publicly available. Experimental results indicate that RiTeK poses significant challenges for current retrieval and LLM systems, while the proposed Relational MCTS method enhances LLM inference ability and achieves state-of-the-art performance on RiTeK.
RiTeK: A Dataset for Large Language Models Complex Reasoning over Textual Knowledge Graphs
Jiatan Huang1thanks: indicates equal contribution, Mingchen Li211footnotemark: 1, Zonghai Yao2, Zhichao Yang2, Yongkang Xiao3 Feiyun Ouyang1, Xiaohan Li4 , Shuo Han1, Hong Yu1,2,5 1University of Massachusetts, Lowell 2University of Massachusetts, Amherst 3University of Minnesota 4Rollins School of Public Health, Emory University 5UMass Chan Medical School 1{Jiatan_Huang, Feiyun_Ouyang, shuo_han,Hong_Yu}@uml.edu 2{mingchenli, zonghaiyao, zhichaoyang}@umass.edu 3xiao0290@umn.edu,4xiaohan.li@emory.edu
1 Introduction
Although large language models (LLMs) have made a significant stride in natural language processing (NLP), complex question answering remains a challenge. Medical professionals, for instance, often need to express complex information that combines flexible inputs with specific, structured constraints. Consider the query: “Which organ or tissue function, that circulates mother and fetus blood, is affected by Fetal Distress?” compared to the simpler version: “What does Fetal Distress affect?” Accurately addressing such complex queries is critical, as it directly impacts healthcare diagnoses and treatment plans.
Figure 1: An example demonstrating the input, expected outputs and system outputs (from GPT-4, GPT-COT, Random Walk, MCTS, Relational MCTS (RMCTS)).
To effectively answer these queries, organizing the underlying knowledge using textual knowledge graphs (TKGs) becomes essential. TKGs integrate unstructured data—such as textual descriptions of nodes (e.g., the definition of the medical term Placental Circulation)—with structured data, like the relationships between entities within the graph (e.g., the relationship between Fetal Distress and Placental Circulation is affects). This integration enables TKGs to represent comprehensive knowledge tailored to specific applications, rendering them invaluable, especially in the medical field, where accuracy and reliability are critically important.
However, existing datasets Wu et al. (2024b, a) exhibit one or more of the following limitations: they are overly simplistic, involving only 1-2 hop reasoning paths; they lack diverse topological structure templates  [^1] A topological structure is a graph that is an abstract of the query graphs of the same pattern, as shown in Li and Ji (2022). and rich relation types; or they fail to incorporate complex constraints  [^2] Constraints are particularly important in KBQA as they help filter out irrelevant information from large knowledge bases, narrowing the search space and improving both efficiency and accuracy. Consequently, these datasets fall short in addressing the complexity of retrieval tasks involving TKGs, especially within the medical domain, where queries require more complex multi-hop reasoning, diverse topological structure templates, and multiple interdependent constraints. Moreover, the absence of textual properties in existing TKGs limits their effectiveness in delivering comprehensive answers.
In addition to the limitations of the TKG datasets themselves, another major challenge lies in how to leverage the useful information embedded in such complex TKGs to enhance the ability of LLMs to answer questions that involve complex reasoning paths within TKGs. As shown in Figure 1, both GPT-4333https://openai.com/index/gpt-4/ and Chain-of-Thought (COT) Wei et al. (2022) fail to generate the correct answer, Nitric oxide. Recent work suggests that utilizing path information within complex KGs can significantly enhance LLM reasoning abilities Li et al. (2024) and provide interpretability. Therefore, exploring accurate and efficient path retrieval with complex topological structures in TKGs is another key focus of the current research community.
To address these gaps, we first present a large-scale complex reasoning dataset over textual knowledge graphs (RiTeK) within the medical domain. In this progress, one primary technical challenge we address is the accurate simulation of user queries with different reasoning types (e.g., six topological structures in Figure 1) within TKGs, ensuring that these queries are relevant and reflective of real-world medical scenarios involving patients, doctors, and medical scientists. This challenge arises from the interdependence of textual and relational information, the complexity of medical terminology and relationships, and the lack of textual descriptions of medical terms. We refer to the framework of Wu et al. (2024b) to simulate user queries and construct precise ground truth answers. However, our focus is primarily on the medical domain, incorporating richer topological structures that extend beyond the traditional 2- and 3-hop structures to better reflect real-world scenarios in the medical domain. Additionally, the textual descriptions of each node are more detailed, enhancing the overall context and understanding. With RiTeK, we delve deeper into retrieval tasks on TKGs, evaluate the capability of current retrieval systems, and provide insights for future advancement.
To address the second challenge and enhance the inference capabilities of LLMs in TKGs within the medical domain, we adapt the Monte Carlo Tree Search Chaslot (2010) to TKGs, and propose an improved Monte Carlo Tree Search, called Relational MCTS to dynamically retrieve relevant relational information for input queries. Our experiments demonstrate that relational MCTS is effective to retrieve the relational path information and enhance the traditional LLM, such as GPT-4.
Our main contributions are summarized as follows. (1) We propose RiTeK, a new QA dataset with complex reasoning over textual knowledge graph in the medical domain. (2) We propose a novel relational MCTS to retrieve the relational information and improve the inference abilities of LLMs within medical TKGs. (3) We delve deeper into retrieval tasks on RiTeK, evaluate the capability of current retrieval systems, and provide insights for future advancement. We also conduct a thorough analysis of our proposed method, including an ablation study, demonstrating the robustness of our framework.
2 Related Work
Datasets of Question Answering over Document. This area of research centers on extracting answers from document sources Rajpurkar (2016); Dunn et al. (2017); Joshi et al. (2017); Trischler et al. (2016); Welbl et al. (2018); Yang et al. (2018); Jin et al. (2021, 2019); Hendrycks et al. (2020). For example, SQuAD Rajpurkar (2016) assesses a model’s ability to interpret and retrieve answers from a single document, focusing on comprehension within a defined context. PubMedQA Jin et al. (2019) focuses on understanding and reasoning within the context of complex biomedical scientific texts. MedQA-CS Yao et al. (2024) aims at simulating authentic clinical scenarios encountered in the clinical examination tasks of medical education. However, unstructured QA datasets often fall short in providing the depth needed for complex relational reasoning to effectively tackle complex user inquiries. In contrast, our research involves queries that demand more complex relational reasoning, challenging the model’s ability to navigate and utilize structured information effectively.
Datasets of Question Answering over Knowledge Graph. The structure QA dataset challenges models to retrieve answers from the structured database, such as knowledge graph Zhang et al. (2018); Yih et al. (2016); Gu et al. (2021); Bao et al. (2016); Trivedi et al. (2017). For example, MetaQA Zhang et al. (2018) challenges models to generate the relational path with multi-hops. To test the models’ abilities to decompose the constraint information in the queries, WebQuestionsSP Yih et al. (2016) is proposed. GrailQA Gu et al. (2021) aims to facilitate the answering of more complex questions, as it allows queries to involve up to four relations and optionally includes functions such as counting, superlatives, and comparatives. However, these datasets primarily focus on relational information, the lack of textual information limits questions within predefined relationships and entities, which constrains the breadth of available information.
Datasets of Question Answering over Textural Knowledge Graph. To integrate textual information into knowledge graphs and queries, STARK (Prime, Amazon, Mag) Wu et al. (2024b) is proposed. To the best of our knowledge, Stark is the only work that focuses on combining relational information with textual information in the question answering over TKGs. However, this dataset lacks sufficient topological structure coverage, hindering the ability to handle complex queries, particularly in the medical domain. The absence of detailed node descriptions further challenges the model’s ability to understand the query information. RiTeK addresses these issues by incorporating richer topological structures and more extensive textual information into knowledge graphs and queries, resulting in more comprehensive and nuanced responses with deeper insights drawn from abundant textual data.
3 Problem Statement
Textural Knowledge Graph
A Textural Knowledge Graph (TKG) is defined as a graph $\mathcal{G}=(\mathcal{R},\mathcal{E},\mathcal{D})$, where $\mathcal{E}$ denotes a set of entities and $\mathcal{R}$ denotes the set of relations among these entities. In a TKG, the entities and relations are usually organized as facts and each fact is defined as a triplet $(h,r,t)$ where $h,t\in\mathcal{E}$ and $r\in\mathcal{R}$ denote the head entity, tail entity and the relation between the two entities, respectively. Each entity $e$ ($e=h$ or $e=t$) in $\mathcal{G}$ has a textual document $d^{e}\in\mathcal{D}$ describing the entity information.
Complex Question Answering over Textural Knowledge Graph
Given a textual knowledge graph $\mathcal{G}$ and input query $q$, the model is expected to generate the answers $a\in\mathcal{E}$, which satisfy the relational constraints defined by the structure of $\mathcal{G}$ as specified in $q$, and the associated document $d^{e}$ need to satisfy the textual requirements in $q$.
Textural Triple Graph
Unlike traditional knowledge graphs, where each node represents an entity and each edge denotes the relationship between nodes, in the textual triple graph, each node corresponds to a triple (head entity, relation, tail entity) along with the textual description of each entity. In this context, the relation indicates whether the two triples are connected. To be specific, let $\mathcal{G*}=(V,E)$ denote a graph consisting of a set of node $V$ and a set of edges $E\in V\times V$. We denote by $n$ the number of nodes in $\mathcal{G}$ and by $m$ its number of edges. Each node $v=(h,r,t,T(h),T(t))\in V$, $T(*)$ is the textual description of entity.
4 Dataset for LLMs Complex Reasoning over TKGs (RiTeK)
Figure 2: The process of constructing textual structured retrieval datasets involves five main steps, 1) Relational template construction: Create the relation template for TKG using the expert-designed topological structure. 2) Extract Textual Properties: Choose one node as the answer node that meets the relational requirement, and extract relevant textual properties. 3) Combine Information: Merge the relational information and textual properties to form a natural-sounding query. 4) Filtering additional answers: Check if the left nodes satisfy the textual properties to establish other ground truth nodes. 5) Expert Evaluation: The medical experts evaluate the naturalness, diversity, and practicality of the dataset.
TKG Dataset # Entities # Relation # Triple # Coverage
Stark-Amazon 4 4 9,443,802
Stark-Mag 4 4 39,802,116
Stark-Prime 10 18 8,100,498 15.29%
RiTeK-PharmKG 3 29 500,958 95.61%
RiTeK-ADint 102 15 1,017,284 36.73%
Table 1: Datasets Statistics of constructed medical textual knowledge graphs. # Coverage refers to textual description coverage of each node. # Entities is the number of entity types. # Relation is the number of relation types. As the textual information for the provided nodes is difficult to analyze statistically, we have not provided the statistical information for Stark Mag and Amazon.
4.1 Medical Textual Knowledge Graph Construction
We construct two medical TKGs based on PharmKG Zheng et al. (2021) and ADint Xiao et al. (2024), as the increased number of entity and relation types introduces significant challenges for path retrieval in the question answering over textual knowledge graph. To enhance the entity attributes, we incorporate textual details from various databases, including Ensembl, UMLS, and Mondo Disease Ontology. As shown in Table 1, our constructed TKGs provide greater node textual coverage, along with more entity types and relation types. For further details on these two medical TKGs, please refer to Appendix A.2.
4.2 Question Answering Dataset Construction
QA Dataset # queries # topological structure # instance rate train/val/test
Stark-Amazon 9,100 1 4 0.65/0.17/0.18
Stark-Mag 13,323 4 1.25 0.60/0.20/0.20
Stark-Prime 1,1204 3 9.3 0.55/0.20/0.25
RiTeK-PharmKG 1,0235 6 11.33 0.80/0.10/0.10
RiTeK-ADint 5322 6 9.67 0.80/0.10/0.10
Table 2: Statistical Overview of the Textual KBQA benchmark Datasets. Instance rate refers to the average number of relational templates per topological structure.
4.2.1 Overview
We developed two question answering datasets RiTeK PharmKG and RiTeK ADint based on two textual knowledge graphs for complex reasoning. These datasets notably feature queries that integrate relational and textual knowledge, incorporating relational templates with broader coverage and higher instance rates. Additionally, to enhance their applicability in practical scenarios, these queries mimic real-life query patterns, exhibiting a natural-sounding quality and flexible formats. Specifically, RiTeK-PharmKG consisting of 10,235 synthesized queries. To maximize the coverage of the question topology, we generate the queries following the six types of topological structure (e.g., multi-hop and multi-hop with constraints). For the synthesized queries, we developed 68 relational templates, crafted by a medical expert and detailed in Appendix C.1, to encompass various relation types and ensure practical relevance. The instance rate of 11.33, which is higher than that of the current TKG dataset Start (Amazon, Mag, and Prime), highlights the higher diversity of this dataset. RiTeK-ADint consisting of 5322 synthesized queries also covering 6 topological structures, with 58 relational templates, and detailed in Appendix C.2. To capture diverse language styles used by different users, same as Stark, we also simulate three distinct roles: medical scientist, doctor, and patient. We divide the synthesized queries on each dataset into training, validation, and testing subsets, with the ratios detailed in Table  2, we use random splits to divide the data. Further details on the scale of our QA benchmarks can be found in Table 2.
4.2.2 Construction Pipline
We present the pipeline that generates the large-scale medical QA datasets on TKGs. The core idea is to entangle relational information and textual properties in to the query, and accurately construct ground truth answers with more topological structures. The construction of the QA datasets (Figure 2) generally involves five steps, and the specific processes vary depending on the characteristics of each dataset. These steps are as follows:
Relational Template Construction. As shown in Figure 2 Step1. We first created templates based on the 6 designed topological structures Li and Ji (2022), and these templates were evaluated by medical experts to ensure their practical relevance and value. Afterward, the topological structures are instantiated conceptually with experts. For instance, for the topological structure Head entity–relation–tail entity, the "(antibiotic) causes <pathologic function>" is a valid and common medical relation template, as antibiotics, particularly penicillin and cephalosporins, are well-known for triggering drug hypersensitivity reactions. This makes it a medically reasonable and frequently observed relationship. We then converted these relation templates into specific relationship queries, such as "Anti-Bacterial Agents causes pathologic function." Since each query could correspond to one or more candidate entities, we matched the queries with the textual KG to obtain k candidate entities.
Extracting Textual Properties. As shown in Figure 2 Step2. For the k candidate answers that meet the relationship criteria, we select one entity as the gold answer and use GPT-4 to extract textual properties from the entity’s associated document. For instance, in the relationship "Anti-Bacterial Agents causes pathologic function," we selected "Hypersensitivity" as the gold answer and extracted textual properties, as shown in Figure 2 (Texutal properties). These textual properties expand upon the concept of hypersensitivity, highlighting its key characteristics, which makes it more likely to fulfill the inquirer’s needs.
Combining Textual and Relational Information. As shown in Figure 2 Step3. After obtaining the relationship templates and textual properties, we combine these components to synthesize the queries. We chose GPT-4 as the LLM for query synthesis, as it excels at generating natural, human-like questions. Additionally, we optimized the prompt and incorporated instructions for different personas to make the queries more diverse and realistic. This approach enhances the quality of our dataset and increases the demands on our model’s reasoning capabilities.
Filtering Additional Answers. As shown in Figure 2 Step4. In addition to the gold answer from which the textual properties are extracted, we need to evaluate whether other candidates meet the requirements of the query in order to include them in the final answer set. We use multiple LLMs to assess whether each candidate’s description satisfies the textual requirements of the query. Only candidates that pass validation by all LLMs will be added to the final answer set.
Human Evaluation. We invited four medical experts to evaluate 100 synthetic queries sampled from two datasets. The evaluation was conducted using a 5-point Likert-like scale across three dimensions. Naturalness refers to how grammatically and logically human-like the queries sound. Diversity assesses whether the queries exhibit complex logical structures and encompass multiple entities, relations, and textual requirements. Practicality measures the real-world applicability of the generated queries and their likelihood of being used in everyday scenarios. The scores were ultimately converted into percentages representing the rates of Positive and Acceptable responses. As shown in the Table 3, our two datasets received high ratings across all three dimensions. The average acceptable rates across the three dimensions were 95.6%, 99.3%, and 82.8%, respectively.
The experts observed that certain queries present issues related to practicality. For instance, the query, What chemical, associated with the metabolism and pharmacokinetics of the gene ’mif’, is also soluble in water? demonstrates a misalignment between the question and the answer. The proposed answer, nitric oxide, lacks sufficient evidence to directly support its involvement in the metabolism of the ’mif’ gene. Furthermore, queries such as What is a disease that is an ancestor of cerebrovascular disorders and can result in brain ischemia? are more suitable for medical education contexts rather than for everyday question-answering scenarios. In contrast, queries like What binding of TF to channels affects the expression/production of NAT1, resulting in the chemical that helps prevent megaloblastic anemia? provide sufficient relational context and textual clarity, making it easier to arrive at the correct answer.
Naturalness Diversity Practicality
RiTeK-PharmKG 77.3/93.3 97.3/99.3 65.3/86.0
RiTeK-ADint 80.2/97.9 95.9/99.3 65.9/79.6
Table 3: Positive/Acceptable rates(%) from experts
4.2.3 Data Distribution Analysis
We chose Shannon Entropy and Type-Token Ratio (TTR) as the metrics to evaluate the query diversity generated in our two datasets. Shannon Entropy takes into account the frequency of each word, measuring the evenness of word distribution in the text, while Type-Token Ratio reflects the variety of words, with a higher value indicating greater diversity in the generated queries. As shown in Table 4, the TTR values for both RiTeK-PharmKG and RiTeK-ADint surpass those of STARK-Prime, demonstrating that the queries generated in our datasets exhibit high complexity and diversity. For Shannon Entropy, our results are comparable to STARK-Prime. Since our RiTeK-ADint dataset involves a wide range of non-pharmacological interventions (NPIs), lifestyle modifications, and environmental factors, it introduces a richer variety of specialized terminology and concepts into the synthesized queries. This expanded vocabulary diversity leads to significantly higher Shannon Entropy compared to the other medical domain datasets. However, since our two datasets are derived from the medical domain, the frequent repetition of specialized medical terminology, as well as the more concentrated vocabulary compared to general-domain texts, results in slightly lower Shannon Entropy for our datasets than for the other two general-domain datasets. For more analysis about the distribution of query lengths and answer length, please refer Appendix A.3.
5 Proposed Approach: Relational MCTS
5.1 Motivation and Approach Overview
Monte Carlo Tree Search (MCTS) has made significant advancements in mathematics, as it dynamically explores and evaluates potential solutions, balancing exploration and exploitation to optimize decision-making in complex, high-dimensional spaces. However, the effectiveness of MCTS in TKGs has not yet been explored. To address this gap, this paper investigates the effectiveness of MCTS and proposes an improved version Relational MCTS that dynamically retrieves relational paths from the TKGs for the special query. Specifically, it involves four steps: 1) To improve the retrieval efficiency, we begin our path retrieval by creating the textual triple graph (TTG) for each textual KG. 2) Subsequently, the node embeddings in the TTG and the input query embedding are computed using a pre-trained language model, such as a SentenceBERT Baek et al. (2023). Notably, each node embedding includes its associated textual information. These embeddings will be used to calculate the reward in the MCTS and Relational MCTS. 3) Proposed Relation MCTS is used to search the relational path information. 4) Finally, the retrieved relational path information will be verbalized and fed into the LLM, such as GPT-4, along with the input query to predict the final answers. Please refer to Appendix B for detailed information on steps 2 and 3.
Shannon Entropy Type-Token Ratio
Medical domain
RiTeK-ADint 10.04 0.187
RiTeK-PharmKG 9.61 0.157
STARK-PRIME 9.63 0.143
General domain
STARK-AMAZON 10.39 0.179
STARK-MAG 10.25 0.180
Table 4: Query diversity measurement
5.2 MCTS vs Relational MCTS
In traditional MCTS, a trajectory typically extends from the root to a terminal node. However, in the context of QA tasks over textual knowledge graphs, many relational templates incorporate constraints, which pose challenges for standard MCTS. Furthermore, the depth of the reasoning path can vary depending on the specific question being posed; for example, some queries may require a reasoning depth of three hops, while others may only require two. Consequently, MCTS is hard to appropriately recognize or adapt to the required depth for each individual query. To address this issue, we propose an enhanced MCTS specifically designed to retrieve the relevant inference path information for the input query, which we refer to as Relational MCTS.
Figure 3: Expansion progress of MCTS and Relational MCTS. The pink shallow circle refers to the $Stop$ sign. Each node $tn$ refers to a triple.
Same as traditional MCTS, starting from the initial topic entity $e$, we perform multiple searches consisting of selection, expansion, simulations, and back-propagation. The key distinction between MCTS and Relational MCTS lies in the expansion phase, as illustrated in Figure 3. In the expansion phase of MCTS, the next move is typically selected from the set of children of the current node. For instance, in MCTS, the candidate moves from node $t1$ are $t3$ and $t4$. In contrast, in Relational MCTS, the candidate moves from node $t1$ include $t3$, $t4$, a $Stop$ action, and the sibling nodes of $t1$. At the same depth level, if two sibling nodes exist, such as in Step 4 of Relational MCTS, the next move is selected from the children of these sibling nodes. For example, when node $t4$ is selected, its candidate moves are $t5$, $t6$, and the $Stop$ action. The inclusion of the $Stop$ action allows Relational MCTS to automatically halt and generate a relevant relational path (1 or 2 hops) for the input query.
RiTeK-PharmKG RiTeK-ADint
Exact Match Rouge-1 Exact Match Rouge-1
Approach P R F1 P R F1 P R F1 P R F1
Zero-Shot GPT-4 11.39 10.90 11.03 15.56 15.50 15.30 7.26 12.10 8.03 13.71 27.64 16.35
+COT Wei et al. (2022) 13.11 16.42 13.70 17.53 22.57 18.40 10.52 19.78 11.95 17.79 37.25 20.97
+Random Walk 12.27 11.86 11.96 14.69 14.15 14.30 15.12 22.68 16.52 20.87 32.92 23.25
+MCTS  Chaslot (2010) 17.17 16.54 16.68 19.09 18.44 18.60 16.97 24.41 18.35 22.82 34.69 25.20
+Relational MCTS (ours) 17.39 20.06 17.57 20.50 23.90 20.82 17.45 24.62 18.76 23.08 34.59 25.27
Few-Shot GPT-4 13.75 15.54 14.04 16.84 19.84 17.49 17.57 17.91 17.48 25.50 28.08 26.04
+COT Wei et al. (2022) 17.29 16.91 16.99 21.55 20.97 21.13 18.57 18.12 18.26 26.68 26.62 26.53
+Random Walk 11.02 13.28 11.32 14.46 17.88 14.92 22.99 22.79 22.75 29.10 29.07 28.95
+MCTS Chaslot (2010) 17.79 17.11 17.30 20.97 20.29 20.48 19.51 27.32 20.91 24.71 36.25 26.96
+Relational MCTS (ours) 19.82 19.03 19.21 22.74 21.81 22.06 23.69 23.72 23.48 30.09 30.36 29.98
Table 5: Results of various approaches for question answering with complex reasoning on RiTeK-PharmKG and RiTeK-ADint
6 Experiments
6.1 Retrieval Models and Evaluation Metrics
We evaluated five different models on our benchmark datasets, including:
  • GPT-4 Achiam et al. (2023): We use GPT-4 with the instruction to generate the answers directly.
  • Chain-of-Thought Wei et al. (2022): We designed the instruction to guide GPT-4 in generating the answer step by step, with the output formatted as step-by-step reasoning: explanation, answer: medical terms.
  • Random Walk: Starting from the topic entity, a random walk algorithm is applied to explore paths in the textual triple graph in the maximum depth $d$.
  • MCTS Chaslot et al. (2008): Starting from the topic entity, a monte carlo tree search algorithm is applied to explore paths in the textual triple graph in the maximum depth $d$.
  • Relational MCTS.
We evaluated the outputs of different methods using several metrics, including Exact Match (EM)Rajpurkar (2016); Li and Huang (2023), which assesses whether the predicted sequence exactly matches the reference, awarding credit only for perfect matches. Additionally, we employed ROUGE-1Cohan and Goharian (2016) to measure unigram overlap between the predicted and reference sequences, providing partial credit for shared words even when the sequences are not identical. To ensure fairness in the comparison, the instructions and examples are the same for both the zero-shot and few-shot settings, respectively.
6.2 Results and Discussion
Table 5 shows the experiment results of various approaches based on Excat Match and Rougle-1. We have the following observations. (1) We observed that the baseline models struggle to generate the correct answers on RiTeK-PharmKG and RiTeK-ADint. For GPT-4 and GPT+COT, they are challenges in utilizing reasoning information from the graph. Although GPT+COT can utilize step-by-step reasoning, it still relies on the inherent knowledge of the LLM, which limits its ability to apply clear logical reasoning based on knowledge graphs. For the Random Walk, while it can provide reasoning paths, its random nature limits its ability to accurately identify the correct path information. However, it could get the better performance than GPT-4 in RiTek-Adint. Both MCTS and Relational MCTS still face challenges in consistently answering questions and providing the correct reasoning path in 100 percent. However, they can offer valuable path information that enhances LLM performance, especially for Relational LLMs. (2) Our Relational MCTS outperforms all the strong baselines across all evaluation metrics. This success is attributed to the designed stop sign and expansion progress, which help the model identify the appropriate path information effectively. (3) Relational MCTS demonstrates better precision but lower recall in the RiTeK-ADint dataset. This discrepancy arises from the fact that MCTS generates an average of 2.70 answers for each sentence, while RMCTS generates an average of only 1.08 answers. The ground truth average length of answers is 1.02. In the medical domain, we aim to generate more accurate answers.
6.3 The Impact of Hyper-Parameters
In our proposed model, there are two parameters controlling the Relational MCTS: (1) $\epsilon\in\{3000,5000,7000,9000,11000\}$ which is the iteration refers to one complete cycle of the algorithm, where a path is selected, expanded, simulated, and updated to improve decision. (2) $d\in{2,3}$ is the maximum depth of path retrieval. We analyze the impact of these two hyper-parameters in Figure 4 and Figure 5. As we can see, when $\epsilon=5000$, relational MCTS selects paths that are more relevant to the input query and consistently provides better performance than the other settings. Otherwise, we noticed that the model achieves similar performance with $\epsilon=5000$ and $\epsilon=9000$ on the RiTeK-ADint dataset. Since more iterations increase computation time, we ultimately select 5000 iterations. Additionally, $d=3$ outperforms the setting of $d=2$, indicating that when $d=2$, relational paths with more than two hops are not retrieved, which negatively impacts the performance of Relational MCTS. Note: in this section, all experiments are in the zero-shot setting.
Figure 4: The extract match F1 performance of Relational MCTS under different iteration settings.
Figure 5: The extract match F1 performance of Relatinoal MCTS under different hop depth setting
6.4 Case Study
Figure 6: A case study on RiTeK
Finally, we highlight the importance of relational path information retrieval ability in achieving good results on our benchmark dataset. In Figure 6, we conduct a case study to compare the performance of Chain-of-Thought, MCTS and Relational MCTS, they are also based on GPT-4. As shown in Figure 6 (green box), without relational path information, Chain-of-Thought failed to generate the right answer Myocardial Reperfusion Injury. Furthermore, MCTS struggles to retrieve the correct path information, resulting in the retrieval of wrong paths that lead the LLM to generate incorrect entities such as Small RNA and MK2206. In contrast, the relational path retrieved by RMCTS could help LLM find the right answer and generate the relevant reasoning information. Specifically, its reasoning explanation information The path then indicates that the interactions of Small RNA lead to dysfunction in organism function which contributes to the pathologic function of damaged myocardium during reperfusion injury is relevant to the final answers. While in COT, the generated explanation is Ultimately, these changes can lead to conditions like heart failure, which is a pathologic function associated with damaged myocardium. Therefore, the final answer is heart failure, which is not relevant to the final answers. In Figure 6 (red box), Relation MCTS and MCTS both failed to get the reasonable path information for the more complex query. They all failed to get the right relational path set by step. This also highlights the challenges present in our benchmark dataset.
7 Conclusion
We present RiTeK, the first dataset specifically designed to evaluate the capability of models in handling complex reasoning over textual knowledge graphs (TKGs). This dataset offers diverse topological structures, relational types, entity types, and queries that integrate relational and textual information, requiring sophisticated reasoning across TKGs. RiTeK also includes rich textual descriptions for each node. To ensure the authenticity and accuracy of the queries, stringent validation was conducted by medical experts. In addition, to enhance the inference capabilities of LLMs, we are pioneering the adaptation of Monte Carlo Tree Search (MCTS) to textual knowledge graphs (TKGs) by introducing Relational MCTS, which retrieves relational path information along with relevant topological structures. Our extensive experiments on RiTeK reveal significant challenges for current models in handling both textual and complex relational information. Overall, RiTeK offers valuable opportunities for advancing research in complex reasoning tasks over textual knowledge graphs, while Relational MCTS provides a valuable way to retrieving complex relational paths from KGs in an unsupervised manner.
8 Limitations
The primary contribution of our work is the construction of two question-answering datasets based on textual knowledge graphs with complex topological structures. This work reveal significant challenges for current models in effectively handling both textual and complex relational information. We also demonstrated the potential of MCTS for complex reasoning path retrieval over textual knowledge graphs and validated the effectiveness of both MCTS and our proposed relational MCTS. However, the performance of relational MCTS in conjunction with other relational information retrieval methods on knowledge graphs remains unexplored, such as He et al. (2024); Baek et al. (2023) In the future, we will adapt these methods based on knowledge graphs to the context of textual knowledge graphs and investigate the effectiveness of these relational information retrieval methods.
Although we employed four medical experts for human evaluation, increasing the number of qualified domain experts would improve the statistical significance and robustness of our findings. Future work should consider expanding the pool of experts and addressing issues of fairness, and potential biases inherent in LLMs.
Appendix A Appendix
A.1 TKG resources
Ensembl  [^4] https://useast.ensembl.org/index.html,   [^5] https://www.nlm.nih.gov/research/umls/index.html, and Mondo Disease Ontology  [^6] https://mondo.monarchinitiative.org/.
A.2 Medical textual knowledge graph construction
We construct two medical TKGs based on PharmKG Zheng et al. (2021) and ADint Xiao et al. (2024), as the increased number of entity and relation types introduces significant challenges for path retrieval in the question answering over textual knowledge graph. We present the statistics of the relational structure in Table 1 and introduce each TKG as follows:
PharmKG Textual Knowledge Graph: We leverage the existing medical knowledge graph PharmKG Zheng et al. (2021) which is a multi-relational, attribute-rich biomedical knowledge graph (KG) constructed from six publicly available databases that provide high-quality structured information. These databases include OMIM, DrugBank, PharmGKB, Therapeutic Target Database (TTD), SIDER, and HumanNet. PharmKG consists of over 500,000 distinct interconnections between genes, drugs, and diseases, encompassing 29 types of relationships within a vocabulary of approximately 8,000 disambiguated entities. To enhance the entity attributes, we incorporate textual details from various databases, including Ensembl, UMLS, and Mondo Disease Ontology, as supplementary data sources.
ADInt Textual Knowledge Graph: ADIntXiao et al. (2024) is a comprehensive knowledge graph (KG) constructed from biomedical literature, focusing on non-pharmacological interventions (NPI) and their associations with Alzheimer’s disease (AD). ADInt includes 162,212 entities spanning 113 UMLS semantic types, which, upon further classification, consist of 25,604 drugs, 16,474 diseases, 46,060 genes and proteins, 2,525 dietary supplements (DS), and 128 complementary and integrative health (CIH) interventions. Moreover, ADInt contains 1,017,284 triples, capturing 15 distinct relation types, offering a rich dataset for exploring the intricate relationships between NPIs and AD. Same as PharmKG, we also incorporate textual details from various databases, including Ensembl, UMLS, and Mondo Disease Ontology, as supplementary data sources.
A.3 Data Analysis of query length and answer length
We analyzed the distribution of query lengths (i.e., the number of words in each query) to assess the complexity of the queries and the amount of information they contain. As shown in the Figure 7, the query lengths range from 5 to 40 words, with approximately 69% and 61% of queries in the two datasets having lengths between 15 and 25 words.
Then, we analyzed the proportion of ground truth answers associated with each query. Generally, the more ground truth answers there are, the less precise the textual requirements in the query tend to be. To increase the difficulty of the question-answering task, we filtered out queries with too many ground truth answers during the dataset creation process, retaining only those with a maximum of three ground truth answers. In both datasets, over 90% of queries have a single ground truth answer, indicating that our queries are enriched with detailed textual information from entity attributes. This introduces more challenges when developing new graph retrieval methods
Figure 7: Distribution of query lengths and answer lengths on RiTeK-ADint and RiTeK-PharmKG datasets
Appendix B Relation MCTS
B.1 Textural Triplet Graph (TTG) and Input Question Encording
We begin our path retrieval approach by generating node embeddings in the TTG and creating a question embedding using a pre-trained language model. To elaborate, for each node $v$ in $\mathcal{G*}$, we utilize a pre-trained LLM, such as SentenceBert, to calculate the embedding of $v$, yielding the presentation $h_{v}$:
$h_{v}=LM(h:r:t:T(h):T(t))$
where, $T(*)$ is the textual description of head entity $h$ or tail entity $t$. The pre-trained LLM also is used to calculate the embedding of input question $q$, yielding the presentation $h_{q}$:
$h_{q}=LM(q)$
B.2 Relational Information Retrieving
Same as traditional MCTS, starting from the initial topic entity $e$, we perform multiple searches consisting of selection, expansion, simulations, and back-propagation. Specifically, In this approach, simulations are conducted using the default rollout policy, and to improve the accuracy of reward estimation, multiple rollouts are carried out. To maintain a balance between exploration and exploitation, we apply the Upper Confidence Bounds for Trees (UCT) method for node selection. The selection process is represented mathematically as:
$UCT(v,a)=\frac{Q(v,a)}{N(v,a)}+c\sqrt{\frac{lnN_{parent}(v)}{N(v,a)}}$
where $N(v,a)$ denotes the number of times node $v$ has been visited under action $a$ in prior iterations, and $N_{parent}(v)$ represents the visit count of the parent node of $v$. $Q(v,a)$ is the estimated reward value and will be updated through back-propagation. For each node, the reward function is computed based on the similarity between the query and a list that includes the current node and the subsequent node.
$Q(v,a)=Sim(h_{q},h_{v}^{1}:h_{v}^{2}:h_{v}^{d})$
where $d$ is the maximum depth of the searching tree. $c$ is a constant that balances exploitation and exploration. Once the search reaches a terminal node, either a terminal state (e.g. Stop Sign) or a predetermined maximum tree depth $d$, we obtain a trajectory from the root to terminal node.
Appendix C Relational Template
C.1 RiTeK-PharmKG
  1. 1.
    Gene -> [Production by cell population] -> Gene
  2. 2.
    Gene -> [Enhance response, or activate, stimulate] -> Gene
  3. 3.
    Gene -> [Relationships involving regulation and pathways] -> Gene
  4. 4.
    Gene -> [Binding, ligand] -> Gene
  5. 5.
    Gene -> [Affects expression/production] -> Gene
  6. 6.
    Gene -> [Gene-Gene] -> Gene
  7. 7.
    Chemical -> [Chemical-Chemical] -> Chemical
  8. 8.
    Disease -> [Ancestors of disease] -> Disease
  9. 9.
    Disease -> [Associations between diseases] -> Disease
  10. 10.
    Gene -> [Interactions] -> Chemical
  11. 11.
    Chemical -> [Interactions] -> Gene
  12. 12.
    Gene -> [Interactions] -> Gene
  13. 13.
    Gene -> [Interactions] -> Disease
  14. 14.
    Gene -> [Drug targets] -> Disease
  15. 15.
    Gene -> [Role in pathogenesis, or promotes progression] -> Disease
  16. 16.
    Gene -> [Mutations affect, or polymorphisms alter risk] -> Disease
  17. 17.
    Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Gene
  18. 18.
    Disease -> [Overexpression in disease] -> Gene
  19. 19.
    Chemical -> [Treatment or therapy] -> Disease
  20. 20.
    Chemical -> [Side effect or adverse event] -> Disease
  21. 21.
    Chemical -> [Inhibits cell growth] -> Disease
  22. 22.
    Chemical -> [Role in pathogenesis] -> Disease
  23. 23.
    Chemical -> [Prevents, suppresses, or alleviates, reduces] -> Disease
  24. 24.
    Disease -> [Biomarkers (progression)] -> Chemical
  25. 25.
    Chemical -> [Agonism, activation, or antagonism, blocking] -> Gene
  26. 26.
    Chemical -> [Binding, ligand] -> Gene
  27. 27.
    Chemical -> [Affects expression/production] -> Gene
  28. 28.
    Chemical -> [Inhibits] -> Gene
  29. 29.
    Gene -> [Transport, channels] -> Chemical
  30. 30.
    Gene -> [Metabolism, pharmacokinetics] -> Chemical
  31. 31.
    Gene -> [Enzyme activity] -> Chemical
  32. 32.
    Gene -> [Enhance response, or activate, stimulate] -> Gene -> [Drug targets] -> Disease
  33. 33.
    Gene -> [Enhance response, or activate, stimulate] -> Gene -> [Role in pathogenesis, or promotes progression] -> Disease
  34. 34.
    Gene -> [Enhance response, or activate, stimulate] -> Gene -> [Mutations affect, or polymorphisms alter risk] -> Disease
  35. 35.
    Gene -> [Relationships involving regulation and pathways] -> Gene -> [Binding, ligand] -> Gene
  36. 36.
    Gene -> [Binding, ligand] -> Gene -> [Affects expression/production] -> Gene
  37. 37.
    Gene -> [Interactions] -> Gene -> [Interactions] -> Chemical
  38. 38.
    Gene -> [Transport, channels] -> Chemical -> [Agonism, activation, or antagonism, blocking] -> Gene
  39. 39.
    Gene -> [Metabolism, pharmacokinetics] -> Chemical -> [Binding, ligand] -> Gene
  40. 40.
    Gene -> [Enhance response, or activate, stimulate] -> Gene -> [Enhance response, or activate, stimulate] -> Gene
  41. 41.
    Gene -> [Interactions] -> Chemical -> [Treatment or therapy] -> Disease
  42. 42.
    Gene -> [Interactions] -> Chemical -> [Side effect or adverse event] -> Disease
  43. 43.
    Gene -> [Interactions] -> Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Gene
  44. 44.
    Chemical -> [Treatment or therapy] -> Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Gene
  45. 45.
    Disease -> [Associations between diseases] -> Disease -> [Ancestors of disease] -> Disease
  46. 46.
    Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Gene
  47. 47.
    Gene -> [Interactions] -> Gene -> [Transport, channels] -> Chemical
  48. 48.
    Gene -> [Metabolism, pharmacokinetics] -> Chemical -> [Binding, ligand] -> Gene
  49. 49.
    Gene -> [Enhance response, or activate, stimulate] -> Gene -> [Drug targets] -> Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Gene
  50. 50.
    Gene -> [Enhance response, or activate, stimulate] -> Gene -> [Mutations affect, or polymorphisms alter risk] -> Disease -> [Overexpression in disease] -> Gene
  51. 51.
    Gene -> [Transport, channels] -> Chemical -> [Agonism, activation, or antagonism, blocking] -> Gene -> [Binding, ligand] -> Chemical
  52. 52.
    Gene -> [Metabolism, pharmacokinetics] -> Chemical -> [Binding, ligand] -> Gene -> [Inhibits] -> Chemical
  53. 53.
    Gene -> [Interactions] -> Chemical -> [Treatment or therapy] -> Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Gene
  54. 54.
    Gene -> [Interactions] -> Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Gene -> [Transport, channels] -> Chemical
  55. 55.
    Gene -> [Role in pathogenesis, or promotes progression] -> Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Gene -> [Metabolism, pharmacokinetics] -> Chemical
  56. 56.
    Chemical -> [Agonism, activation, or antagonism, blocking] -> Gene -> [Drug targets] -> Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Gene
  57. 57.
    Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Gene -> [Role in pathogenesis, or promotes progression] -> Disease
  58. 58.
    Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Gene -> [Metabolism, pharmacokinetics] -> Chemical -> [Side effect or adverse event] -> Disease
  59. 59.
    Gene -> [Production by cell population] -> Gene -> [Enhance response, or activate, stimulate] -> Gene -> [Relationships involving regulation and pathways] -> Gene
  60. 60.
    Gene -> [Enhance response, or activate, stimulate] -> Gene -> [Binding, ligand] -> Gene -> [Affects expression/production] -> Gene
  61. 61.
    Gene -> [Relationships involving regulation and pathways] -> Gene -> [Gene-Gene] -> Gene -> [Binding, ligand] -> Gene
  62. 62.
    Gene -> [Interactions] -> Gene -> [Interactions] -> Gene -> [Transport, channels] -> Chemical
  63. 63.
    Gene -> [Interactions] -> Gene -> [Interactions] -> Gene -> [Metabolism, pharmacokinetics] -> Chemical
  64. 64.
    Gene -> [Enhance response, or activate, stimulate] -> Gene -> [Mutations affect, or polymorphisms alter risk] -> Disease -> [Overexpression in disease] -> Gene
  65. 65.
    Gene -> [Enzyme activity] -> Chemical -> [Affects expression/production] -> Gene -> [Chemical-Chemical] -> Chemical
  66. 66.
    Gene -> [Interactions] -> Chemical -> [Role in pathogenesis] -> Disease -> [Overexpression in disease] -> Gene
  67. 67.
    Chemical -> [Side effect or adverse event] -> Disease -> [Biomarkers (diagnostic), or regulation linked to disease] -> Gene -> [Mutations affect, or polymorphisms alter risk] -> Disease
  68. 68.
    Chemical -> [Inhibits cell growth] -> Disease -> [Overexpression in disease] -> Gene -> [Role in pathogenesis, or promotes progression] -> Disease
C.2 RiTeK-ADint
  1. 1.
    Amino Acid, Peptide, or Protein -> [affects] -> Cell Function
  2. 2.
    Amino Acid, Peptide, or Protein -> [affects] -> Disease or Syndrome
  3. 3.
    Amino Acid, Peptide, or Protein -> [causes] -> Anatomical Abnormality
  4. 4.
    Amino Acid, Peptide, or Protein -> [interacts with] -> Pharmacologic Substance
  5. 5.
    Anatomical Abnormality -> [affects] -> Organ or Tissue Function
  6. 6.
    Anatomical Abnormality -> [complicates] -> Disease or Syndrome
  7. 7.
    Anatomical Abnormality -> [manifestation of] -> Genetic Function
  8. 8.
    Antibiotic -> [affects] -> Molecular Function
  9. 9.
    Antibiotic -> [causes] -> Pathologic Function
  10. 10.
    Antibiotic -> [disrupts] -> Cell Component
  11. 11.
    Antibiotic -> [treats] -> Disease or Syndrome
  12. 12.
    Bacterium -> [causes] -> Cell or Molecular Dysfunction
  13. 13.
    Bacterium -> [interacts with] -> Human
  14. 14.
    Biologically Active Substance -> [affects] -> Organism Function
  15. 15.
    Biologically Active Substance -> [causes] -> Injury or Poisoning
  16. 16.
    Biologically Active Substance -> [disrupts] -> Gene or Genome
  17. 17.
    Body Part, Organ, or Organ Component -> [produces] -> Immunologic Factor
  18. 18.
    Cell Component -> [affects] -> Molecular Function
  19. 19.
    Cell Component -> [produces] -> Nucleic Acid, Nucleoside, or Nucleotide
  20. 20.
    Cell Function -> [affects] -> Mental or Behavioral Dysfunction
  21. 21.
    Cell Function -> [produces] -> Biologically Active Substance
  22. 22.
    Cell or Molecular Dysfunction -> [affects] -> Neoplastic Process
  23. 23.
    Cell or Molecular Dysfunction -> [manifestation of] -> Pathologic Function
  24. 24.
    Cell -> [produces] -> Organic Chemical
  25. 25.
    Congenital Abnormality -> [affects] -> Virus
  26. 26.
    Congenital Abnormality -> [manifestation of] -> Organism Function
  27. 27.
    Diagnostic Procedure -> [affects] -> Genetic Function
  28. 28.
    Disease or Syndrome -> [affects] -> Organ or Tissue Function
  29. 29.
    Disease or Syndrome -> [associated with] -> Therapeutic or Preventive Procedure
  30. 30.
    Disease or Syndrome -> [manifestation of] -> Cell or Molecular Dysfunction
  31. 31.
    Finding -> [manifestation of] -> Pathologic Function
  32. 32.
    Gene or Genome -> [produces] -> Amino Acid, Peptide, or Protein
  33. 33.
    Genetic Function -> [affects] -> Human
  34. 34.
    Genetic Function -> [produces] -> Cell Component
  35. 35.
    Hazardous or Poisonous Substance -> [affects] -> Mental or Behavioral Dysfunction
  36. 36.
    Hazardous or Poisonous Substance -> [disrupts] -> Organ or Tissue Function
  37. 37.
    Health Care Activity -> [affects] -> Disease or Syndrome
  38. 38.
    Human -> [interacts with] -> Human
  39. 39.
    Immunologic Factor -> [affects] -> Pathologic Function
  40. 40.
    Indicator, Reagent, or Diagnostic Aid -> [interacts with] -> Hazardous or Poisonous Substance
  41. 41.
    Injury or Poisoning -> [disrupts] -> Genetic Function
  42. 42.
    Medical Device -> [treats] -> Mental or Behavioral Dysfunction
  43. 43.
    Mental or Behavioral Dysfunction -> [affects] -> Organism Function
  44. 44.
    Molecular Function -> [affects] -> Virus
  45. 45.
    Neoplastic Process -> [affects] -> Bacterium
  46. 46.
    Neoplastic Process -> [associated with] -> Neoplastic Process
  47. 47.
    Nucleic Acid, Nucleoside, or Nucleotide -> [interacts with] -> Immunologic Factor
  48. 48.
    Organ or Tissue Function -> [produces] -> Immunologic Factor
  49. 49.
    Organic Chemical -> [affects] -> Pathologic Function
  50. 50.
    Organic Chemical -> [interacts with] -> Pharmacologic Substance
  51. 51.
    Organism Function -> [affects] -> Disease or Syndrome
  52. 52.
    Pathologic Function -> [associated with] -> Therapeutic or Preventive Procedure
  53. 53.
    Pathologic Function -> [manifestation of] -> Organ or Tissue Function
  54. 54.
    Pharmacologic Substance -> [affects] -> Genetic Function
  55. 55.
    Pharmacologic Substance -> [treats] -> Sign or Symptom
  56. 56.
    Sign or Symptom -> [manifestation of] -> Genetic Function
  57. 57.
    Therapeutic or Preventive Procedure -> [affects] -> Neoplastic Process
  58. 58.
    Virus -> [interacts with] -> Human